OpenStudy (amberlykhan):

The area of a rectangle is represented by the polynomial x^2 + 3x - 6x - 18. Find the possible expressions for the length and width of the rectangle.

OpenStudy (welshfella):

x^2 + 3x - 6x - 18 = x^2 - 3x - 18 you now need to factor the above

OpenStudy (welshfella):

oh - silly me its easier to factor in its original form you can factor that by grouping the first 2 and last 2 terms

OpenStudy (amberlykhan):

How do you find the length and width expressions after that?

OpenStudy (welshfella):

x^2 + 3x - 6x - 18 = x(x + 3) - 6x - 18 can you find the factors of the last 2 terms?

OpenStudy (amberlykhan):

6.

OpenStudy (amberlykhan):

So it would be -6(x - 3).

OpenStudy (welshfella):

Hint;- its -6 and another expression in paremteses

OpenStudy (amberlykhan):

+3

OpenStudy (welshfella):

right -6(x+ 3) now we have x(x +3) - 6(x + 3) can you continue?

OpenStudy (amberlykhan):

I don't know what to do.

OpenStudy (welshfella):

you can simplify this further Note that (x +3) is common to the 2 'halves' of the expression

OpenStudy (welshfella):

sorry but i have to go right now

OpenStudy (amberlykhan):

Oh, I got it. It would be written as (x - 6)(x + 3), which represents the length x the width. Thanks!